# Trans-dimensional Hamiltonian model selection and parameter estimation from sparse, noisy data

### v1

@@ 1 Introduction
Current approaches to estimating the parameters of the Hamiltonian describing the interaction between a central spin and neighboring nuclear spins, such as correlation spectroscopy, offer high-resolution insights but require long acquisition times and extensive manual tuning [15, 16], making them infeasible at scale. Recently proposed machine learning methods offer some automation [17, 18], but rely on experimental datasets that are expensive to obtain and on a large amount of computing resources. Hence, a framework that can operate in the sparse, noisy data regime and still extract physically meaningful Hamiltonian parameters is desirable. Furthermore, since the number of nuclear spins interacting with the spin-defect is unknown a priori and is itself a quantity to be estimated, the inverse problem at hand is a model dimension selection problem in addition to a Hamiltonian parameter estimation problem.
@@
To address these challenges simultaneously, we present a hybridized Markov chain Monte Carlo (MCMC) framework that combines several specialized techniques to address separately the mathematical challenges that arise. Discrete and continuous random-walk Metropolis-Hastings (RWMH) samplers are used to explore parameter spaces with both continuous and discrete parameters. Trans-dimensional model selection is handled through the reversible-jump MCMC (RJMCMC) algorithm [23], which allows for sampling over models with differing numbers of parameters, a necessity in problems where the model dimension (e.g., the number of nuclear spins) is itself unknown. To better sample from complex and multimodal posteriors in high-dimensional spaces, we incorporate parallel tempering (PT) [24], which improves convergence and avoids trapping in local minima. Together, these methods form a generalized, flexible, and extensible framework to solve inverse problems that can operate robustly in the data-scarce regime. We demonstrate that our hybridized MCMC framework produces posterior distributions that describe simulated nuclear spin baths in the sparse, noisy data regime, and we validate our approach with experimental data to demonstrate that it achieves an accuracy comparable to much more computationally and experimentally intensive procedures, including direct machine learning approaches,[17, 18], but using an order of magnitude fewer data points.
@@ 2.1 General formulation
Furthermore, we acknowledge that “sparse" and “noisy" are quantities that are context-dependent, and we generally consider data $\mathbf{d}_{d}$ to be “sparse" and noisy if there are multiple parameterizations $k_{i},k_{j}$ and $(\mathbf{a}_{k_{i}},\mathbf{b}_{i}),(\mathbf{a}_{k_{j}},\mathbf{b}_{j})$ such that
$\min_{k_{i},\mathbf{a}_{k_{i}}}||f(\mathcal{H}_{k_{i}}(\mathbf{a}_{k_{i}}),\mathbf{b}_{i},\mathbf{c})-\mathbf{d}_{d}||_{N(\mathbf{\epsilon}_{d})}$ $\approx\min_{k_{j},\mathbf{a}_{k_{j}}}||f(\mathcal{H}_{k_{j}}(\mathbf{a}_{k_{j}}),\mathbf{b}_{j},\mathbf{c})-\mathbf{d}_{d}||_{N(\mathbf{\epsilon}_{d})}$ (2)
@@ 3.1 Random walk Metropolis-Hastings (RWMH) for continuous and discrete domains
1:Input: data $\mathbf{d}_{d}$, partition of a fixed-length set of parameters to update $\mathbf{p}=\{p_{1},\dots,p_{m}\}$ and fixed parameters $\mathbf{q}=\{q_{1},\dots,q_{s}\}$, with $\mathbf{p}\cup\mathbf{q}=\{a_{1},\dots,a_{k},b_{1},\dots,b_{j}\}$ and $\mathbf{p}\cap\mathbf{q}=\emptyset$ and their initial values $\mathbf{p}^{(0)}=(p_{1}^{(0)},p_{2}^{(0)},\dots,p_{m}^{(0)})$ and $\mathbf{q}^{(0)}=(q_{1}^{(0)},q_{2}^{(0)},\dots,q_{s}^{(0)})$, known parameters $\mathbf{c}$, fixed Hamiltonian model $\mathcal{H}_{k}(\mathbf{a}_{k})$ indexed by $k$, forward model $f_{k}(\mathcal{H}_{k}(\mathbf{a}_{k}),\mathbf{b},\mathbf{c})=f_{k}(\mathbf{p},\mathbf{q},\mathbf{c})$, likelihood model $\mathcal{L}(\mathbf{d}_{d}|f_{k}(\mathbf{p},\mathbf{q},\mathbf{c})$, domain $\mathcal{D}$, proposal kernel $r_{R,\mathcal{D}}(x,z)$, radius $R$, number of steps $N_{\text{step}}$
2: for $n=0,\dots N_{\text{step}}-1$ do for $n=0,\dots N_{\text{step}}-1$ do
3:      for $n=0,\dots N_{\text{step}}-1$ do for $l=1,\dots,m$ do for $l=1,\dots,m$ do
4:          for $l=1,\dots,m$ do Proposal step: Sample $p^{*}_{l}\sim r_{R,\mathcal{D}}(p_{l}^{(n)},\cdot)$; $\mathbf{p}^{*}=\{p^{(n)}_{1},\dots,p^{(n)}_{l-1},p^{*}_{l},p^{(n)}_{l+1},\dots,p^{(n)}_{m}\}$ Proposal step: Sample $p^{*}_{l}\sim r_{R,\mathcal{D}}(p_{l}^{(n)},\cdot)$; $\mathbf{p}^{*}=\{p^{(n)}_{1},\dots,p^{(n)}_{l-1},p^{*}_{l},p^{(n)}_{l+1},\dots,p^{(n)}_{m}\}$
5:          Proposal step: Sample $p^{*}_{l}\sim r_{R,\mathcal{D}}(p_{l}^{(n)},\cdot)$; $\mathbf{p}^{*}=\{p^{(n)}_{1},\dots,p^{(n)}_{l-1},p^{*}_{l},p^{(n)}_{l+1},\dots,p^{(n)}_{m}\}$ Accept/reject step: Update $\mathbf{p}^{(n+1)}$ =$=\begin{cases}\mathbf{p}^{*}&\text{with probability }\alpha(\mathbf{p}^{(n)},\mathbf{p}^{*})=\min\{1,\frac{\mathcal{L}(\mathbf{d}_{d}|f_{k}(\mathbf{p}^{*},\mathbf{q},\mathbf{c})}{\mathcal{L}(\mathbf{d}_{d}|f_{k}(\mathbf{p}^{(n)},\mathbf{q},\mathbf{c})}\frac{r_{R,\mathcal{D}}(p^{*}_{l},p^{(n)}_{l})}{r_{R,\mathcal{D}}(p^{(n)}_{l},p^{*}_{l})}\}\\ \mathbf{p}^{(n)}&\text{with probability }1-\alpha(\mathbf{p}^{(n)},\mathbf{p}^{*})\end{cases}$ $\mathbf{q}^{(n+1)}$ $=\mathbf{q}^{(n)}$ Accept/reject step: Update $\mathbf{p}^{(n+1)}$ =$=\begin{cases}\mathbf{p}^{*}&\text{with probability }\alpha(\mathbf{p}^{(n)},\mathbf{p}^{*})=\min\{1,\frac{\mathcal{L}(\mathbf{d}_{d}|f_{k}(\mathbf{p}^{*},\mathbf{q},\mathbf{c})}{\mathcal{L}(\mathbf{d}_{d}|f_{k}(\mathbf{p}^{(n)},\mathbf{q},\mathbf{c})}\frac{r_{R,\mathcal{D}}(p^{*}_{l},p^{(n)}_{l})}{r_{R,\mathcal{D}}(p^{(n)}_{l},p^{*}_{l})}\}\\ \mathbf{p}^{(n)}&\text{with probability }1-\alpha(\mathbf{p}^{(n)},\mathbf{p}^{*})\end{cases}$ $\mathbf{q}^{(n+1)}$ $=\mathbf{q}^{(n)}$ Accept/reject step: Update $\mathbf{p}^{(n+1)}$ =$=\begin{cases}\mathbf{p}^{*}&\text{with probability }\alpha(\mathbf{p}^{(n)},\mathbf{p}^{*})=\min\{1,\frac{\mathcal{L}(\mathbf{d}_{d}|f_{k}(\mathbf{p}^{*},\mathbf{q},\mathbf{c})}{\mathcal{L}(\mathbf{d}_{d}|f_{k}(\mathbf{p}^{(n)},\mathbf{q},\mathbf{c})}\frac{r_{R,\mathcal{D}}(p^{*}_{l},p^{(n)}_{l})}{r_{R,\mathcal{D}}(p^{(n)}_{l},p^{*}_{l})}\}\\ \mathbf{p}^{(n)}&\text{with probability }1-\alpha(\mathbf{p}^{(n)},\mathbf{p}^{*})\end{cases}$ $\mathbf{q}^{(n+1)}$ $=\mathbf{q}^{(n)}$
@@ 3.2 Reverse Jump Markov Chain Monte Carlo (RJMCMC)
2: for $n=0,\dots N_{\text{step}}-1$ do for $n=0,\dots N_{\text{step}}-1$ do
3:      for $n=0,\dots N_{\text{step}}-1$ do Proposal step: sample $k^{*}\sim\gamma(k^{(n)},\cdot)$ Proposal step: sample $k^{*}\sim\gamma(k^{(n)},\cdot)$
4:      Proposal step: sample $k^{*}\sim\gamma(k^{(n)},\cdot)$ if $k^{*}>k^{(n)}$ then $k^{*}$ $=k^{(n)}+1$ $p^{*}_{k+1}$ $\sim\mathcal{D}$ $\mathbf{p}^{*}$ $=(p_{1}^{(n)},p_{2}^{(n)},\dots,p_{k}^{(n)},p_{k+1}^{*})$ if $k^{*}>k^{(n)}$ then $k^{*}$ $=k^{(n)}+1$ $p^{*}_{k+1}$ $\sim\mathcal{D}$ $\mathbf{p}^{*}$ $=(p_{1}^{(n)},p_{2}^{(n)},\dots,p_{k}^{(n)},p_{k+1}^{*})$
5:      if $k^{*}>k^{(n)}$ then $k^{*}$ $=k^{(n)}+1$ $p^{*}_{k+1}$ $\sim\mathcal{D}$ $\mathbf{p}^{*}$ $=(p_{1}^{(n)},p_{2}^{(n)},\dots,p_{k}^{(n)},p_{k+1}^{*})$ else $k^{*}$ $=k^{(n)}-1$ $l$$\sim\{1,2,\dots,k^{(n)}\}$ $\mathbf{p}^{*}$ $=(p_{1}^{(n)},\dots,p_{l-1}^{(n)},p_{l+1}^{(n)},\dots,p_{k}^{(n)})$ else $k^{*}$ $=k^{(n)}-1$ $l$$\sim\{1,2,\dots,k^{(n)}\}$ $\mathbf{p}^{*}$ $=(p_{1}^{(n)},\dots,p_{l-1}^{(n)},p_{l+1}^{(n)},\dots,p_{k}^{(n)})$ else $k^{*}$ $=k^{(n)}-1$ $l$$\sim\{1,2,\dots,k^{(n)}\}$ $\mathbf{p}^{*}$ $=(p_{1}^{(n)},\dots,p_{l-1}^{(n)},p_{l+1}^{(n)},\dots,p_{k}^{(n)})$
6:      else $k^{*}$ $=k^{(n)}-1$ $l$$\sim\{1,2,\dots,k^{(n)}\}$ $\mathbf{p}^{*}$ $=(p_{1}^{(n)},\dots,p_{l-1}^{(n)},p_{l+1}^{(n)},\dots,p_{k}^{(n)})$ Accept/reject step: Update $k^{n+1},\mathbf{p}^{(n+1)}$ =$=\begin{cases}k^{*},\mathbf{p}^{*}&\text{with probability }\alpha(k^{(n)},\mathbf{p}^{(n)};k^{*},\mathbf{p}^{*})\\ k^{(n)},\mathbf{p}^{(n)}&\text{with probability }1-\alpha(k^{(n)},\mathbf{p}^{(n)};k^{*},\mathbf{p}^{*})\end{cases}$ $\mathbf{q}^{(n+1)}$ $=\mathbf{q}^{(n)}$ $\alpha(k^{(n)},\mathbf{p}^{(n)};k^{*},\mathbf{p}^{*})$ $=\min\{1,\frac{\mathcal{L}(\mathbf{d}_{d}|f_{k^{*}}(\mathbf{p}^{*},\mathbf{q},\mathbf{c})}{\mathcal{L}(\mathbf{d}_{d}|f_{k^{(n)}}(\mathbf{p}^{(n)},\mathbf{q},\mathbf{c})}\frac{\gamma(k^{*},k^{(n)})}{\gamma(k^{(n)},k^{*})}\}$ Accept/reject step: Update $k^{n+1},\mathbf{p}^{(n+1)}$ =$=\begin{cases}k^{*},\mathbf{p}^{*}&\text{with probability }\alpha(k^{(n)},\mathbf{p}^{(n)};k^{*},\mathbf{p}^{*})\\ k^{(n)},\mathbf{p}^{(n)}&\text{with probability }1-\alpha(k^{(n)},\mathbf{p}^{(n)};k^{*},\mathbf{p}^{*})\end{cases}$ $\mathbf{q}^{(n+1)}$ $=\mathbf{q}^{(n)}$ $\alpha(k^{(n)},\mathbf{p}^{(n)};k^{*},\mathbf{p}^{*})$ $=\min\{1,\frac{\mathcal{L}(\mathbf{d}_{d}|f_{k^{*}}(\mathbf{p}^{*},\mathbf{q},\mathbf{c})}{\mathcal{L}(\mathbf{d}_{d}|f_{k^{(n)}}(\mathbf{p}^{(n)},\mathbf{q},\mathbf{c})}\frac{\gamma(k^{*},k^{(n)})}{\gamma(k^{(n)},k^{*})}\}$ Accept/reject step: Update $k^{n+1},\mathbf{p}^{(n+1)}$ =$=\begin{cases}k^{*},\mathbf{p}^{*}&\text{with probability }\alpha(k^{(n)},\mathbf{p}^{(n)};k^{*},\mathbf{p}^{*})\\ k^{(n)},\mathbf{p}^{(n)}&\text{with probability }1-\alpha(k^{(n)},\mathbf{p}^{(n)};k^{*},\mathbf{p}^{*})\end{cases}$ $\mathbf{q}^{(n+1)}$ $=\mathbf{q}^{(n)}$ $\alpha(k^{(n)},\mathbf{p}^{(n)};k^{*},\mathbf{p}^{*})$ $=\min\{1,\frac{\mathcal{L}(\mathbf{d}_{d}|f_{k^{*}}(\mathbf{p}^{*},\mathbf{q},\mathbf{c})}{\mathcal{L}(\mathbf{d}_{d}|f_{k^{(n)}}(\mathbf{p}^{(n)},\mathbf{q},\mathbf{c})}\frac{\gamma(k^{*},k^{(n)})}{\gamma(k^{(n)},k^{*})}\}$
@@ 3.3 Parallel Tempering (PT)
2: for $n=0,\dots N_{\text{step}}-1$ do for $n=0,\dots N_{\text{step}}-1$ do
@@
3:      for $n=0,\dots N_{\text{step}}-1$ do Within-strand step: Generate $\tilde{\mathbf{p}}_{j}^{(n+1)}$ from $\mathbf{p}_{j}^{(n+1)}$ by one step of (Alg. 1) for each strand $j\in[1,\dots,J]$ using the corresponding likelihood model $\mathcal{L}_{j}(\mathbf{d}_{d}|f_{k}(\mathbf{p},\mathbf{q},\mathbf{c}))=(\mathcal{L}(\mathbf{d}_{d}|f_{k}(\mathbf{p},\mathbf{q},\mathbf{c})))^{\beta_{j}}$ Within-strand step: Generate $\tilde{\mathbf{p}}_{j}^{(n+1)}$ from $\mathbf{p}_{j}^{(n+1)}$ by one step of (Alg. 1) for each strand $j\in[1,\dots,J]$ using the corresponding likelihood model $\mathcal{L}_{j}(\mathbf{d}_{d}|f_{k}(\mathbf{p},\mathbf{q},\mathbf{c}))=(\mathcal{L}(\mathbf{d}_{d}|f_{k}(\mathbf{p},\mathbf{q},\mathbf{c})))^{\beta_{j}}$
4:      Within-strand step: Generate $\tilde{\mathbf{p}}_{j}^{(n+1)}$ from $\mathbf{p}_{j}^{(n+1)}$ by one step of (Alg. 1) for each strand $j\in[1,\dots,J]$ using the corresponding likelihood model $\mathcal{L}_{j}(\mathbf{d}_{d}|f_{k}(\mathbf{p},\mathbf{q},\mathbf{c}))=(\mathcal{L}(\mathbf{d}_{d}|f_{k}(\mathbf{p},\mathbf{q},\mathbf{c})))^{\beta_{j}}$ Draw strands to attempt swap: Sample $a,b\sim\{1,\dots J\}$ with $a\neq b$ uniformly at random Draw strands to attempt swap: Sample $a,b\sim\{1,\dots J\}$ with $a\neq b$ uniformly at random
@@
5:      Draw strands to attempt swap: Sample $a,b\sim\{1,\dots J\}$ with $a\neq b$ uniformly at random Evolve non-swapped chains forward: For $j\notin\{a,b\}$, set $\mathbf{p}_{j}^{(n+1)},\mathbf{q}_{j}^{(n+1)}=\tilde{\mathbf{p}}_{j}^{(n+1)},\mathbf{q}_{j}^{(n)}$ Evolve non-swapped chains forward: For $j\notin\{a,b\}$, set $\mathbf{p}_{j}^{(n+1)},\mathbf{q}_{j}^{(n+1)}=\tilde{\mathbf{p}}_{j}^{(n+1)},\mathbf{q}_{j}^{(n)}$
@@
6:      Evolve non-swapped chains forward: For $j\notin\{a,b\}$, set $\mathbf{p}_{j}^{(n+1)},\mathbf{q}_{j}^{(n+1)}=\tilde{\mathbf{p}}_{j}^{(n+1)},\mathbf{q}_{j}^{(n)}$ Swap proposal step: Attempt a swap of states between strands $a$ and $b$: $(\mathbf{p}_{a}^{(n+1)},\mathbf{p}_{b}^{(n+1)})$ =$=\begin{cases}(\tilde{\mathbf{p}}_{b}^{(n+1)},\tilde{\mathbf{p}}_{a}^{(n+1)})&\text{with probability }\alpha_{\text{PT}}(a,b)\\ (\tilde{\mathbf{p}}_{a}^{(n+1)},\tilde{\mathbf{p}}_{b}^{(n+1)})&\text{with probability }1-\alpha_{\text{PT}}(a,b)\end{cases}$ $\mathbf{q}_{a}^{(n+1)},\mathbf{q}_{b}^{(n+1)}$ $=\mathbf{q}_{a}^{(n)},\mathbf{q}_{b}^{(n)}$ $\alpha_{\text{PT}}$ $=\min\{1,\frac{\mathcal{L}_{a}(\mathbf{d}_{d}|\tilde{\mathbf{p}}_{b}^{(n+1},\mathbf{q},\mathbf{c})}{\mathcal{L}_{a}(\mathbf{d}_{d}|\tilde{\mathbf{p}}_{a}^{(n+1},\mathbf{q},\mathbf{c})}\frac{\mathcal{L}_{b}(\mathbf{d}_{d}|\tilde{\mathbf{p}}_{a}^{(n+1)},\mathbf{q},\mathbf{c})}{\mathcal{L}_{b}(\mathbf{d}_{d}|\tilde{\mathbf{p}}_{b}^{(n+1)},\mathbf{q},\mathbf{c})}\}$ Swap proposal step: Attempt a swap of states between strands $a$ and $b$: $(\mathbf{p}_{a}^{(n+1)},\mathbf{p}_{b}^{(n+1)})$ =$=\begin{cases}(\tilde{\mathbf{p}}_{b}^{(n+1)},\tilde{\mathbf{p}}_{a}^{(n+1)})&\text{with probability }\alpha_{\text{PT}}(a,b)\\ (\tilde{\mathbf{p}}_{a}^{(n+1)},\tilde{\mathbf{p}}_{b}^{(n+1)})&\text{with probability }1-\alpha_{\text{PT}}(a,b)\end{cases}$ $\mathbf{q}_{a}^{(n+1)},\mathbf{q}_{b}^{(n+1)}$ $=\mathbf{q}_{a}^{(n)},\mathbf{q}_{b}^{(n)}$ $\alpha_{\text{PT}}$ $=\min\{1,\frac{\mathcal{L}_{a}(\mathbf{d}_{d}|\tilde{\mathbf{p}}_{b}^{(n+1},\mathbf{q},\mathbf{c})}{\mathcal{L}_{a}(\mathbf{d}_{d}|\tilde{\mathbf{p}}_{a}^{(n+1},\mathbf{q},\mathbf{c})}\frac{\mathcal{L}_{b}(\mathbf{d}_{d}|\tilde{\mathbf{p}}_{a}^{(n+1)},\mathbf{q},\mathbf{c})}{\mathcal{L}_{b}(\mathbf{d}_{d}|\tilde{\mathbf{p}}_{b}^{(n+1)},\mathbf{q},\mathbf{c})}\}$ Swap proposal step: Attempt a swap of states between strands $a$ and $b$: $(\mathbf{p}_{a}^{(n+1)},\mathbf{p}_{b}^{(n+1)})$ =$=\begin{cases}(\tilde{\mathbf{p}}_{b}^{(n+1)},\tilde{\mathbf{p}}_{a}^{(n+1)})&\text{with probability }\alpha_{\text{PT}}(a,b)\\ (\tilde{\mathbf{p}}_{a}^{(n+1)},\tilde{\mathbf{p}}_{b}^{(n+1)})&\text{with probability }1-\alpha_{\text{PT}}(a,b)\end{cases}$ $\mathbf{q}_{a}^{(n+1)},\mathbf{q}_{b}^{(n+1)}$ $=\mathbf{q}_{a}^{(n)},\mathbf{q}_{b}^{(n)}$ $\alpha_{\text{PT}}$ $=\min\{1,\frac{\mathcal{L}_{a}(\mathbf{d}_{d}|\tilde{\mathbf{p}}_{b}^{(n+1},\mathbf{q},\mathbf{c})}{\mathcal{L}_{a}(\mathbf{d}_{d}|\tilde{\mathbf{p}}_{a}^{(n+1},\mathbf{q},\mathbf{c})}\frac{\mathcal{L}_{b}(\mathbf{d}_{d}|\tilde{\mathbf{p}}_{a}^{(n+1)},\mathbf{q},\mathbf{c})}{\mathcal{L}_{b}(\mathbf{d}_{d}|\tilde{\mathbf{p}}_{b}^{(n+1)},\mathbf{q},\mathbf{c})}\}$

### v2

@@ 1 Introduction
Current approaches to estimating the parameters of the Hamiltonian describing the interaction between a central spin and neighboring nuclear spins, such as correlation spectroscopy, offer high-resolution insights but require long acquisition times and extensive manual tuning [15, 16], making them infeasible at scale. Recently proposed machine learning methods offer some automation [17, 18], but rely on experimental datasets that are expensive to obtain and rely on a large amount of computing resources. Hence, a framework that can operate in the sparse, noisy data regime and still extract physically meaningful Hamiltonian parameters is desirable. Furthermore, since the number of nuclear spins interacting with the spin-defect is unknown a priori and is itself a quantity to be estimated, the inverse problem at hand is a model dimension selection problem in addition to a Hamiltonian parameter estimation problem.
@@
To address these challenges simultaneously, we present a hybridized Markov chain Monte Carlo (MCMC) framework that combines several specialized techniques to address separately the mathematical challenges that arise. Discrete and continuous random-walk Metropolis-Hastings (RWMH) samplers are used to explore parameter spaces with both continuous and discrete parameters. Trans-dimensional model selection is handled through the reversible-jump MCMC (RJMCMC) algorithm [23], which allows for sampling over models with differing numbers of parameters, a necessity in problems where the model dimension (e.g., the number of nuclear spins) is itself unknown. To better sample from complex and multimodal posteriors in high-dimensional spaces, we incorporate parallel tempering (PT) [24], which improves convergence and avoids trapping in local minima. Together, these methods form a generalized, flexible, and extensible framework to solve inverse problems that can operate robustly in the data-scarce regime. We demonstrate that our hybridized MCMC framework produces posterior distributions that describe simulated nuclear spin baths in the sparse, noisy data regime, and we validate our approach with experimental data to demonstrate that it achieves an accuracy comparable to much more computationally and experimentally intensive procedures, including direct machine learning approaches [17, 18], but using an order of magnitude fewer data points.
@@ 2.1 General formulation
Furthermore, we acknowledge that “sparse" and “noisy" are quantities that are context-dependent, and we generally consider data $\mathbf{d}_{d}$ to be “sparse" and “noisy" if there are multiple parameterizations $k_{i},k_{j}$ and $(\mathbf{a}_{k_{i}},\mathbf{b}_{i}),(\mathbf{a}_{k_{j}},\mathbf{b}_{j})$ such that
[…]
@@ 3.1 Random walk Metropolis-Hastings (RWMH) for continuous and discrete domains
1:Input: Data $\mathbf{d}_{d}$, partition of a fixed-length set of parameters to update $\mathbf{p}=\{p_{1},\dots,p_{m}\}$ and fixed parameters $\mathbf{q}=\{q_{1},\dots,q_{s}\}$, with $\mathbf{p}\cup\mathbf{q}=\{a_{1},\dots,a_{k},b_{1},\dots,b_{j}\}$ and $\mathbf{p}\cap\mathbf{q}=\emptyset$ and their initial values $\mathbf{p}^{(0)}=(p_{1}^{(0)},p_{2}^{(0)},\dots,p_{m}^{(0)})$ and $\mathbf{q}^{(0)}=(q_{1}^{(0)},q_{2}^{(0)},\dots,q_{s}^{(0)})$, known parameters $\mathbf{c}$, fixed Hamiltonian model $\mathcal{H}_{k}(\mathbf{a}_{k})$ indexed by $k$, forward model $f_{k}(\mathcal{H}_{k}(\mathbf{a}_{k}),\mathbf{b},\mathbf{c})=f_{k}(\mathbf{p},\mathbf{q},\mathbf{c})$, likelihood model $\mathcal{L}(\mathbf{d}_{d}|f_{k}(\mathbf{p},\mathbf{q},\mathbf{c}))$, domain $\mathcal{D}$, proposal kernel $r_{R,\mathcal{D}}(x,z)$, radius $R$, number of steps $N_{\text{step}}$
2: for $n=0,\dots N_{\text{step}}-1$ do
3:   for $l=1,\dots,m$ do
4:    Proposal step: Sample $p^{*}_{l}\sim r_{R,\mathcal{D}}(p_{l}^{(n)},\cdot)$; $\mathbf{p}^{*}=\{p^{(n)}_{1},\dots,p^{(n)}_{l-1},p^{*}_{l},p^{(n)}_{l+1},\dots,p^{(n)}_{m}\}$
5:    Accept/reject step: Update $\mathbf{p}^{(n+1)}$ $=\begin{cases}\mathbf{p}^{*}&\text{with probability }\alpha(\mathbf{p}^{(n)},\mathbf{p}^{*})=\min\{1,\frac{\mathcal{L}(\mathbf{d}_{d}|f_{k}(\mathbf{p}^{*},\mathbf{q},\mathbf{c}))}{\mathcal{L}(\mathbf{d}_{d}|f_{k}(\mathbf{p}^{(n)},\mathbf{q},\mathbf{c}))}\frac{r_{R,\mathcal{D}}(p^{*}_{l},p^{(n)}_{l})}{r_{R,\mathcal{D}}(p^{(n)}_{l},p^{*}_{l})}\}\\ \mathbf{p}^{(n)}&\text{with probability }1-\alpha(\mathbf{p}^{(n)},\mathbf{p}^{*})\end{cases}$ $\mathbf{q}^{(n+1)}$ $=\mathbf{q}^{(n)}$
@@ 3.2 Reverse Jump Markov Chain Monte Carlo (RJMCMC)
2: for $n=0,\dots N_{\text{step}}-1$ do
3:   Proposal step: sample $k^{*}\sim\gamma(k^{(n)},\cdot)$
4:   if $k^{*}>k^{(n)}$ then $k^{*}$ $=k^{(n)}+1$ $p^{*}_{k+1}$ $\sim\mathcal{D}$ $\mathbf{p}^{*}$ $=(p_{1}^{(n)},p_{2}^{(n)},\dots,p_{k}^{(n)},p_{k+1}^{*})$
5:   else $k^{*}$ $=k^{(n)}-1$ $l$ $\sim\{1,2,\dots,k^{(n)}\}$ $\mathbf{p}^{*}$ $=(p_{1}^{(n)},\dots,p_{l-1}^{(n)},p_{l+1}^{(n)},\dots,p_{k}^{(n)})$
6:   Accept/reject step: Update $k^{n+1},\mathbf{p}^{(n+1)}$ $=\begin{cases}k^{*},\mathbf{p}^{*}&\text{with probability }\alpha(k^{(n)},\mathbf{p}^{(n)};k^{*},\mathbf{p}^{*})\\ k^{(n)},\mathbf{p}^{(n)}&\text{with probability }1-\alpha(k^{(n)},\mathbf{p}^{(n)};k^{*},\mathbf{p}^{*})\end{cases}$ $\mathbf{q}^{(n+1)}$ $=\mathbf{q}^{(n)}$ $\alpha(k^{(n)},\mathbf{p}^{(n)};k^{*},\mathbf{p}^{*})$ $=\min\{1,\frac{\mathcal{L}(\mathbf{d}_{d}|f_{k^{*}}(\mathbf{p}^{*},\mathbf{q},\mathbf{c})}{\mathcal{L}(\mathbf{d}_{d}|f_{k^{(n)}}(\mathbf{p}^{(n)},\mathbf{q},\mathbf{c})}\frac{\gamma(k^{*},k^{(n)})}{\gamma(k^{(n)},k^{*})}\}$
@@ 3.3 Parallel Tempering (PT)
2: for $n=0,\dots N_{\text{step}}-1$ do
@@
3:   Within-strand step: Generate $\tilde{\mathbf{p}}_{j}^{(n+1)}$ from $\mathbf{p}_{j}^{(n+1)}$ by one step of (Alg. 1) for each strand $j\in[1,\dots,J]$ using the corresponding likelihood model $\mathcal{L}_{j}(\mathbf{d}_{d}|f_{k}(\mathbf{p},\mathbf{q},\mathbf{c}))=(\mathcal{L}(\mathbf{d}_{d}|f_{k}(\mathbf{p},\mathbf{q},\mathbf{c})))^{\beta_{j}}$
@@
4:   Draw strands to attempt swap: Sample $a,b\sim\{1,\dots J\}$ with $a\neq b$ uniformly at random
5:   Evolve non-swapped chains forward: For $j\notin\{a,b\}$, set $\mathbf{p}_{j}^{(n+1)},\mathbf{q}_{j}^{(n+1)}=\tilde{\mathbf{p}}_{j}^{(n+1)},\mathbf{q}_{j}^{(n)}$
@@
6:   Swap proposal step: Attempt a swap of states between strands $a$ and $b$: $(\mathbf{p}_{a}^{(n+1)},\mathbf{p}_{b}^{(n+1)})$ $=\begin{cases}(\tilde{\mathbf{p}}_{b}^{(n+1)},\tilde{\mathbf{p}}_{a}^{(n+1)})&\text{with probability }\alpha_{\text{PT}}(a,b)\\ (\tilde{\mathbf{p}}_{a}^{(n+1)},\tilde{\mathbf{p}}_{b}^{(n+1)})&\text{with probability }1-\alpha_{\text{PT}}(a,b)\end{cases}$ $\mathbf{q}_{a}^{(n+1)},\mathbf{q}_{b}^{(n+1)}$ $=\mathbf{q}_{a}^{(n)},\mathbf{q}_{b}^{(n)}$ $\alpha_{\text{PT}}$ $=\min\{1,\frac{\mathcal{L}_{a}(\mathbf{d}_{d}|\tilde{\mathbf{p}}_{b}^{(n+1},\mathbf{q},\mathbf{c})}{\mathcal{L}_{a}(\mathbf{d}_{d}|\tilde{\mathbf{p}}_{a}^{(n+1},\mathbf{q},\mathbf{c})}\frac{\mathcal{L}_{b}(\mathbf{d}_{d}|\tilde{\mathbf{p}}_{a}^{(n+1)},\mathbf{q},\mathbf{c})}{\mathcal{L}_{b}(\mathbf{d}_{d}|\tilde{\mathbf{p}}_{b}^{(n+1)},\mathbf{q},\mathbf{c})}\}$